what is...
theoretical probability
experimental probability
fraction operations
probability with replacement
and punnett squares

Answer by stanbon(75887)

what is...
theoretical probability
The teoretical prob. of getting a 3 when you roll a die is 1/6;
but if you actually roll a die 1000 times and find that the 3
comes up half of the time, the experimental probability of
getting a three is 1/2.
experimental probability
----------------------------
fraction operations
Sounds like arithmetic to me.
------------------------------
probability with replacement
If you have a jar with 3 red and 2 blue marbles the probability
of selecting a red and then a blue depends on whether you
"replace" the 1st drawn marble.
With replacement the probability is (3/5)(2/5) = 6/25
without replacement the prob is (3/5)(2/4) = 6/20
--------
"With replacement" means you are conducting the same experiment
on each draw. Without replacement means the experiment is
changing on each draw.
----------
With replacement means the different draw results are independent.
Without replacement means the different draw results are dependent.
===============
and punnett squares
I have no idea.
================
Cheers,
Stan H.